OpenStudy (anonymous):

Evaluate. ArcTan( sqrt3/ 3)

OpenStudy (anonymous):

\[\tan^{-1} \frac{ \sqrt{3} }{ 3 }\]

geerky42 (geerky42):

Let \(\theta\) be \( \tan^{-1} \dfrac{\sqrt{3}}{3} \). \(\tan \theta = \dfrac{\sqrt{3}}{3}\) At what angle is \(\tan \theta\) equal to \( \dfrac{\sqrt3}{3} \)?

OpenStudy (anonymous):

It's 30. But I don't know how I got that.

OpenStudy (anonymous):



geerky42 (geerky42):

Well, \(\tan 30^{\text{o}} = \dfrac{\sqrt3}{3}\) so it's true that \(30^{\text{o}} = \tan^{-1}\dfrac{\sqrt3}{3}\).

OpenStudy (anonymous):

I thought Tangent was opposite over adjacent.

OpenStudy (anonymous):

OH THEY RATIONALIZED IT

geerky42 (geerky42):

It is. Welcome.

OpenStudy (anonymous):

God damn it, tricky bastards. >:| Thanks a lot. xD